A Chemist needs 120 milliliters of a 21% solution but only has 3% and 27% solutions available. Find how many milliliters of 27% solution should be used to the desired solution.

Answer by Alan3354(23885)

A Chemist needs 120 milliliters of a 21% solution but only has 3% and 27% solutions available. Find how many milliliters of 27% solution should be used to the desired solution.
-----------------
t = amount of 3%
s = amount of 27%
----
t + s = 120
3t + 27s = 120*21
----
Can you do the rest?